the number of atoms in 20g of SO3 is approximately
  1. 1×1023
  2. 1.5×1023
  3. 2×1023
  4. 6×1023

Grade:12th pass

4 Answers

Arun
4 years ago
Molecular mass of SO3 --- 32+3x16=80 Unified mass

Molar Mass=80gm
Given weight of SO3=20g
Number moles=given weight/gram molecular weight=20/80=1/4=0.25 moles
So 1 mole =80 gm of SO3 contains 6.023x 10²³ moelcules
0.25 moles[20g] conatins -----?
=6.023x10²³x0.25
=1.50x10²³ moleucles
Each moleucle of SO3 has 4 atoms.
so,Total number of atoms in SO3 is 4x1.50x10²³ 
=6x10²³ atoms
∴Total number of atoms in SO3 is 6x10²³ atoms
 
Saurabh Koranglekar
4 years ago
Dear student

Moles present in 20 g of SO3 = 20/ 80.07 = 0.2498 mol.
​Number of molecules = 0.2498 X 6.022 X 1023 = 1.5 X 1023 molecules.
Each mole of SO3 has 6.022 X 1023 molecules.

Regards
Khimraj
4 years ago

79 g of SO3 contains 1 mole of molecules or 4 mole atoms(3 mole atoms of oxygen and 1 mole atoms of sulphur).

 

So,Number of atoms in20 g SO3

=4×6.02×10^23×20÷79 = 6×10^23

Answer=6 × 10^23

Answerian
3 years ago
 
Molecular mass of SO3  = 32+3x16=80 Unified mass
Molar Mass = 80g
Given weight of SO3 = 20g
Number moles = given weight/gram molecular weight
                            =20/80=1/4
                            =0.25 moles
So;
1 mole = 80 g of SO3 contains 6.023 x 10²³ moelcules
0.25 moles[20g] conatins = ?
                         = 6.023x10²³ x 0.25
                         = 1.50x10²³ moleucles
Each moleucle of SO3 has 4 atoms.
So,
Total number of SO3 atoms are = 4 x 1.50x10²³ 
                                                         =6x10²³ atoms
∴Total number of atoms in SO3 are = 6x10²³ atoms